The lengths of the sides of a triangle are consecutive odd integers. If the perimeter is 1 less than four times the shortest side, find the lengths of each side.
Answer by user_dude2008(1862)

x+(x+2)+(x+4)=4x-1

3x+6=4x-1

3x-4x=-1-6

-x=-7

x=7

Lengths: 7, 9, 11